# Deep Learning Based Dynamics Identification and Linearization of Orbital Problems using Koopman Theory

### v2

@@ 5.3 Training Loss
In this section we analyze the training loss for both models. We can see that from both Figures 16 and 17, the loss function exponentially decays as the number of training epochs increases. This is very typical behaviour of a NN model, and shows that the model is adequately trained for each prediction due to its steady state value reached. With more training, tuning of the hyperparameters and loss function it could be reduced even further.
Figure 16: Neural Network training loss across all training epochs for the 2BP.
Figure 17: Neural Network training loss across all training epochs for the CR3BP.
@@ 5.4 Discussion
The propagation of the 2BP in the Earth system is very accurate, with minimal error in both the position states and metrics, even for IC’s that were outside of the trained region. The largest absolute error in all of the trialed IC’s in the Earth system was 3 km, which for that orbit translates to an error percentage of 0.01%. This percentage is calculated as the ratio of the largest error over the mean radius of the orbit.
@@
By utilizing the LTI Koopman operator learned on the Earth system on the Moon and Jupiter systems, we see that there is a capability for the model to generalize to other dynamical systems, but the model is sensitive to variations in parameters in the nonlinear dynamics. The error in the predictions varied in comparison to the Earth model because of the dissimilar mass of the central body, ultimately changing its gravitational parameter. The largest error percentage in the Moon system was $\sim 0.008\%$ whilst the largest error in the Jupiter system was $\sim 0.02\%$. Although the errors are still very small, if such a system was to be used in a controls application or a trajectory tracking situation, these errors would have to be reduced further, or have the control system correct it.
@@
The analysis of the accuracy of the 2BP Koopman model with the metrics in Section 4.1 shows that our model not only is able to predict the current trajectory of the orbit, but also is able to conserve the physical properties that are evident in the nonlinear dynamics. The learned linear model is even able to conserve these invariants for systems it has never seen before or been trained on, emphasising the remarkable capabilities of the Koopman operator to effectively linearize a nonlinear system, whilst remaining physically and mathematically valid. Likewise the analysis of the Jacobi Constant for the CR3BP model highlights its ability to conserve the physical properties related with its dynamics. Although the accuracy of the CR3BP LTI model is not as high as the 2BP model, it clearly illustrates the ability to capture key important features, with the relative error in the Jacobi Constant being $\lesssim 12.7\%$. Both the 2BP and CR3BP cases show that the model is accurately learning the dynamics in the linearization and not simply learning shapes or curves.
@@ 6 CONCLUSION
The advancements in this paper are significant because, to our knowledge, this is the first time that both the 2BP and CR3BP have been globally linearized. Our proposed model, which is also capable of generalization to a wide array of celestial systems, demonstrates a fast and accurate method for developing a linear Koopman operator representation of the 2BP and CR3BP. We establish the accuracy of both models through analysis of the invariant properties of each nonlinear system, and show that they hold for our linearized model.
@@
7 FUTURE WORK
Investigations of the capacity to generalize the Koopman operator to various orbits and resolve the effects of the gravitational parameter on its predictions is an immediate area of improvement in this model. Furthermore, the work with the CR3BP model suggests that the Koopman operator is reasonably capable of learning complex dynamics. Hence, a realistic two-body model, such as one including drag, solar radiation pressure and a more precise gravitational potential of the Earth (e.g., J2 gravitational parameter), would allow for a real-world application of this Koopman operator.
The hyperparameters of the DNN in this work could be further explored, as many of them were empirically chosen and adjust until the final parameters were chosen. Increasing the number of neurons and hidden layers may allow for improved prediction. An investigation into whether it is more important to have a higher latent space or higher number of hidden layers or both, would be relevant in addressing the larger error of the CR3BP and how that aspect may be mitigated.
For implementing practical applications of the framework, the Koopman model should be expanded with EDMDc [27] (i.e., EDMD with control) in order to learn the control matrix required for a real-world control application of the linearized dynamics. It is worth noting that accurate state and control data should be paired when learning the model based on EDMDc.

### v3

@@ 5.3 Training Loss
In this section we analyze the training loss for both models. We can see that from both Figures 28(a) and 28(b), the loss function exponentially decays as the number of training epochs increases. This is very typical behavior of a NN model, and shows that the model is adequately trained for each prediction due to its steady state value reached. With more training, tuning of the hyperparameters and loss function it could be reduced even further.
(a) Neural Network training loss across all training epochs for the 2BP.
(b) Neural Network training loss across all training epochs for the CR3BP.
@@ 5.4 Discussion
The propagation of the 2BP in the Earth system is very accurate, with minimal global and local error in both the position states and metrics, even for IC’s that were outside of the trained region. The largest absolute error in all of the trialed IC’s in the Earth system was 2 km, which for that orbit translates to an error percentage of 0.018% (5000 km orbit). This percentage is calculated as the ratio of the largest error over the mean radius of the orbit. A longer time evolution for the prediction was considered in the next simulation, where 10 orbital periods where prediction from each initial condition. The maximum error percentage experienced in this simulation was $0.18\%$ (5000 km orbit), which although higher than the single orbit, is still relatively low. We notice however that this error grows steadily as the length of the prediction increases, so methods of state estimation in a complete guidance navigation and control loop with this LTI Koopman model would likely help alleviate this growing error by ’resetting’ the initial condition.
@@
By utilizing the LTI Koopman operator learned on the Earth system on the Moon and Jupiter systems, we see that there is a capability for the model to generalize to other dynamical systems, however, the model is sensitive to variations in parameters in the nonlinear dynamics. The error in the predictions varied in comparison to the Earth model because of the dissimilar mass of the central body, ultimately changing its gravitational parameter. The largest error percentage in the Moon system was $\sim 0.0033\%$ (30,000 km orbit) whilst the largest error in the Jupiter system was $\sim 0.015\%$ (30,000 km orbit). Although the errors are still very small, if such a system was to be used in a controls application or a trajectory tracking situation, these errors would have to be reduced further, or have the control system correct it.
We next demonstrate the capability of the model to handle nonlinear and non-periodic disturbances to the nonlinear dynamics in the form of the J2 gravity perturbation and the solar radiation pressure force. We notice adequate agreeance between the exact nonlinear dynamics and the LTI Koopman model, with a maximum error percentage of $0.14\%$ (300 km orbit). The ability of the model to generalize to eccentric orbits showed that this was where the model struggled with the most. The larger the eccentricity of the orbit, the worse the prediction became, with an eccentricity of 0.5 displaying noticeable inconsistencies between the linear and nonlinear systems. The positional error percentage for these orbits however had a maximum value of $0.04\%$ (300 km orbit). With further concentration in terms of training data, loss function manipulation and hyperparameter tuning it would be possible to tune the model to provide higher degrees of accuracy for a specific application that requires eccentric or perturbed orbits.
The analysis of the accuracy of the 2BP Koopman model with the metrics in Section 4.2 shows that our model not only is able to predict the current trajectory of the orbit, but also is able to conserve the physical properties that are evident in the nonlinear dynamics. The learned linear model is even able to conserve these invariants for systems it has never seen before or been trained on, emphasizing the remarkable capabilities of the Koopman operator to effectively linearize a nonlinear system, whilst remaining physically and mathematically valid. Deviating from the purely circular orbits, it is evident that the model struggles with preserving the $r\cdot v$ metric for some of the more complex models simulated. The 300 km orbit around the Moon, 5000 km orbit for ten periods, and a majority of the perturbed circular orbits experienced undesirable values in this metric. Our implementation of the metric-specific loss function reduced the magnitude of the disparity by an order of two, showing it is possible to reduce this further, however conclusive investigation is needed.
@@
The analysis of the Jacobi Constant for the CR3BP model highlights its ability to conserve the physical properties related with its dynamics. Although the accuracy of the CR3BP LTI model is not as high as the 2BP model, it clearly illustrates the ability to capture key important features, with the relative error in the Jacobi Constant being $\lesssim 12.7\%$. Both the 2BP and CR3BP cases show that the model is accurately learning the dynamics in the linearization and not simply learning shapes or curves.
@@ 6 CONCLUSION
The advancements in this paper are significant because, to our knowledge, this is the first time that circular, eccentric and perturbed 2BP and the CR3BP, around the L1 Lagrange point, have been globally linearized. Our proposed model, which is also capable of generalization to a wide array of celestial systems, demonstrates a fast and accurate method for developing a linear Koopman operator representation of the 2BP and CR3BP. We establish the accuracy of both models through analysis of the invariant properties of each nonlinear system, and show that each of the invariants hold for our linearized model.